OpenStudy (anonymous):

Inverse Variation Written Assignment..In need of help with all of it, if you can please help...don't have much time.

OpenStudy (anonymous):

What is the assignment?

OpenStudy (anonymous):

btw, inverse variation is just y=k/x

OpenStudy (whpalmer4):

What don't you understand about the graphing? Pick an x value. Calculate the y value. Make a mark at that spot. Repeat.

OpenStudy (anonymous):

just plot each point and then draw a curve through all the points that you graphed to best fit it

OpenStudy (anonymous):

So would I graph the data from A or B or both? I don't understand...I don't even know what its supposed to look like to be perfectly honest. lol

OpenStudy (anonymous):

i would make 2 separate graphs

OpenStudy (whpalmer4):

Two graphs. Here's a graph of an inverse variation:

OpenStudy (whpalmer4):

when one value gets large, the other gets small, and vice versa.

OpenStudy (anonymous):

Okay so do I graph these numbers exactly how they are...like the weight side would be x and the distance would be y or what...?

OpenStudy (anonymous):

im sorry I've just never been shown how to do these right..

OpenStudy (whpalmer4):

Yes, graph weight along the x axis and distance along the y axis. Usually, the independent variable goes on the x axis and the dependent variable goes on the y axis.

OpenStudy (anonymous):

So I wouldn't need to make these numbers smaller? I'm just trying to picture in my head what this is supposed to look like...cause so far I got nothing. I was looking at this page http://math.tutorvista.com/number-system/inverse-variation.html but all of these numbers are big and I've never graphed from a table..ugh the life of home school...

OpenStudy (whpalmer4):

Look, just make your tick marks be bigger chunks of number. Instead of 1, 2, 3, make them 10,20,30, or 100, 200, 300, or whatever gives you a reasonably sized graph.

OpenStudy (whpalmer4):

Also, there's nothing that says your graph has to go from 0 to 3000 if all of the action in the graph is between 1800 and 2200. Just start the graph at 1700 and go to 2300 in that case.

OpenStudy (whpalmer4):

totally made up numbers, only for illustration of the point, not related to your homework!

OpenStudy (anonymous):

The first graph I did has a straight line when I plotted the points, it looks like a direct variation graph...is that okay or did I do something wrong?

OpenStudy (whpalmer4):

It shouldn't be a completely straight line, but it doesn't bend much for those points, either. Here's what I did. x's mark the data points.

OpenStudy (whpalmer4):

and for the second one:

OpenStudy (anonymous):

Ohhh..do the small numbers go on the bottom of the graph?? i wonder if i have my x and y axis confused ...my second one looks like yours though :)

OpenStudy (whpalmer4):

The data in the first graph is showing \[y = \frac{90000}{x}\]I plotted both the data from the table (marked with the x's) and then the function for a much wider range, so you can see the data you have in context:

OpenStudy (anonymous):

Oh okay...how'd you get the y to equal 9000?? cause i have 900 and 100 in the table...I'm starting to get something though so thank you for helping :]]

OpenStudy (whpalmer4):

the relationship is \[y = \frac{k}{x}\]If we know that \(y=900\) when \(x = 100\) (or vice versa), then \[900 = \frac{k}{100}\]\[100*900 = 100*\frac{k}{100}\]\[90000 = k\]\[y = \frac{90000}{x}\]

OpenStudy (whpalmer4):

Notice that in that last graph, the y-axis is shown at x = 400, NOT x = 0...

OpenStudy (whpalmer4):

I should have adjusted that for you, here it is in the usual spot:

OpenStudy (anonymous):

Oh okay!! thank you...=) alright so is that good for the graphs?? Because then it says Does it look as though there is a relationship between the weight and its distance from the fulcrum and what type of variation is there? I'm curious what is the fulcrum...

OpenStudy (anonymous):

Oh well I gotta go, but will definitely be back for more help tomorrow.!

OpenStudy (whpalmer4):

the fulcrum is the pivot on the arm...

OpenStudy (whpalmer4):

the spot where the seesaw is held up.